Fermi National Laboratory


Accelerator Update

The integrated luminosity for last week was 11.7 pb-1.

Monday March 1
The day shift began with Operations monitoring store 3263, the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

At 8:39 AM, a Linac quadrupole power supply (QPS104) started to drift. Technicians worked on the supply and had it running properly by 9:47 AM.

A Booster expert began a collimator study at 10:48 AM.

At 12:43 PM, A Booster RF station (BRF13) tripped off for the first of nine times over a three-hour period.

Pbar began a reverse protons study at 1:24 PM. This study halted stacking.

At 1:44 PM, Main Injector (MI) and Recycler experts began conducting studies.

Tevatron (TeV) experts began a study period around 10 PM.

Pbar studies ended and stacking resumed at 10:49 PM.

A TeV RF station (TRF5) tripped off at 10:55 PM due to its water heaters tripping off. An expert said not to reset the heaters until after the TeV finished its studies.

At 11:35 PM, Operations reported odd things happening with the MI BPMs. BPM and frontend reboots didn't help. Upon further investigation, operators discovered it was a timing problem, which they fixed.

Tuesday March 2
The midnight shift began with Operations monitoring store 3263, with Pbar stacking, and with studies in TeV and MI.

At 2:25 AM, the TeV experts completed their studies and Operations terminated store 3263.

Operators tried to reset the TeV TRF5 water heaters at 2:46 AM, but they tripped off. Operators called in experts.

Operations put the TeV through a dry squeeze at 2:47 AM.

The RF experts replaced a power supply for the TRF5 water heater controller. They had the station up and running by 3:36 AM.

Operations put the TeV into shot setup at 5:15 AM.

At 7:16 AM, the TeV suffered a sector C1 quench while ramping. Pbar resumed stacking.

Operators had the Collider Detector at Fermilab (CDF) and D-zero (D0) collision halls prepared for access by 8 AM.

Both CDF and D0 were out of access by 4 PM.

At 5:30 PM, Operations put the TeV through a dry squeeze.

Operations put the TeV into shot setup at 8:44 PM.

At 9:42 PM, a Linac RF station (LRF4) tripped off. Operators investigated and found a LCW leak. A Linac tech came in and repaired a hose. He had the station operating by 11:10 PM. Shot setup resumed.

Wednesday March 3
The midnight shift began with the TeV in shot setup and with MiniBooNE conducting a study.

At 12:35 AM, Operations established store 3271 with an initial luminosity of 52.03E30. Pbar resumed stacking.

A Linac vacuum pump for LRF4 tripped at 2:45 AM. It reset okay.

At 6 AM, Operations and Recycler experts began stashing antiprotons from Pbar. By 7:45 AM, Pbar had sent eight shots of antiprotons to the Recycler.

Today's Plans
The plans for today include completing the Recycler stashes, monitoring store 3271, and stacking.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 3/3/2004   email Fermilab